# Professionally translated parallel data for 115 under-represented languages

### v1

@@ Appendix F Full results
gaa | Both | +23.1 | 22.7 | 44.5 | 43.7 | 45.8 | 47.4 | 34.7 | 44.0 | 40.9 | 6.6 | 48.3 | - |
dov | Both | +21.1 | 19.1 | 39.2 | 39.4 | 40.2 | 40.6 | 19.2 | 39.5 | 18.2 | 8.7 | 41.7 | - |
ahr | SmolDoc | +17.8 | 24.2 | 31.8 | 41.5 | 42.0 | 42.8 | 32.8 | 39.0 | 30.0 | 36.9 | - | - |
sus | Both | +17.8 | 11.3 | 28.3 | 27.0 | 29.1 | 30.3 | 26.1 | 29.4 | 20.7 | 5.6 | 34.6 | - |
nqo | Both | +17.5 | 0.2 | 17.9 | 17.3 | 17.7 | 17.5 | 17.1 | 17.9 | 17.2 | 1.1 | 19.1 | - |
alz | Both | +15.5 | 16.9 | 31.5 | 31.7 | 32.4 | 33.4 | 25.3 | 30.6 | 26.9 | 8.9 | 36.6 | - |
lu | Both | +13.8 | 27.6 | 37.5 | 39.0 | 41.4 | 42.2 | 27.2 | 37.0 | 34.8 | 21.9 | - | - |
@@
ks | SmolSent | +7.3 | 22.1 | 29.4 | 0.4 | 29.4 | 29.7 | 28.0 | 30.4 | 30.5 | 26.3 | - | 36.7 |
nr | Both | +7.0 | 48.9 | 54.0 | 50.2 | 55.9 | 57.5 | 48.0 | 53.6 | 51.2 | 45.5 | 59.5 | - |
doi | SmolDoc | +6.6 | 34.3 | 28.1 | 40.8 | 40.9 | 41.3 | 35.9 | 39.5 | 38.2 | 27.7 | 40.4 | - |
sat-L. | SmolSent | +6.4 | 12.8 | 19.5 | 15.9 | 19.2 | 20.8 | 25.3 | 22.5 | 22.7 | 21.3 | 22.7 | - |
mfe | Both | +5.3 | 59.5 | 65.4 | 61.7 | 64.8 | 66.9 | 59.6 | 65.0 | 59.8 | 59.5 | 67.5 | - |
ach | Both | +5.2 | 33.2 | 43.1 | 32.3 | 38.4 | 39.2 | 32.4 | 37.3 | 35.1 | 23.8 | 43.2 | - |
ayl | Both | +4.5 | 47.3 | 51.7 | 51.3 | 51.8 | 53.9 | 45.3 | 48.9 | 46.3 | 48.6 | - | - |
ber-L. | Both | +4.2 | 26.1 | 27.9 | 29.7 | 30.3 | 30.9 | 27.6 | 32.7 | 32.1 | 21.2 | 34.7 | - |
apd-S. | Both | +3.6 | 42.3 | 50.2 | 43.8 | 45.9 | 47.1 | 43.2 | 45.0 | 42.5 | 45.6 | - | - |
ve | Both | +3.5 | 47.9 | 50.0 | 49.5 | 51.4 | 52.2 | 50.2 | 53.1 | 52.7 | 43.9 | 56.8 | - |
kri | Both | +2.8 | 31.5 | 34.2 | 32.1 | 34.3 | 34.7 | 34.5 | 33.5 | 30.7 | 34.9 | 34.9 | - |
tiv | Both | +2.5 | 23.8 | 25.7 | 25.5 | 26.3 | 26.5 | 22.3 | 24.2 | 24.5 | 1.5 | 25.2 | - |
gn | SmolSent | +2.3 | 37.4 | 37.8 | 29.4 | 39.7 | 38.4 | 36.0 | 38.0 | 36.4 | 35.6 | 38.4 | 38.5 |
mos | Both | +2.3 | 18.2 | 20.9 | 19.2 | 20.5 | 21.1 | 24.3 | 25.0 | 20.9 | 1.3 | - | 23.8 |
ar-M. | Both | +0.1 | 40.1 | 41.8 | 38.7 | 40.2 | 40.9 | 40.5 | 40.8 | 40.4 | 41.0 | - | 43.0 |
arz | Both | -0.1 | 48.6 | 46.1 | 48.6 | 48.5 | 47.8 | 49.6 | 50.3 | 48.8 | 49.7 | - | 49.6 |
kl | SmolSent | -0.3 | 40.6 | 39.7 | 29.0 | 40.3 | 41.2 | 42.2 | 43.1 | 41.2 | 41.6 | 42.9 | - |
sa | SmolSent | -0.3 | 33.0 | 32.9 | 27.4 | 32.7 | 33.4 | 31.8 | 33.0 | 32.1 | 32.0 | 35.2 | 29.0 |
ay | SmolSent | -0.5 | 32.7 | 31.8 | 23.2 | 32.2 | 32.4 | 33.4 | 33.2 | 32.9 | 30.0 | 34.7 | 31.7 |
@@
bci | Both | -0.7 | 23.2 | 22.2 | 22.4 | 22.5 | 22.9 | 17.1 | 20.7 | 27.6 | 1.0 | 29.8 | - |
ndc-Z. | Both | -1.0 | 29.2 | 27.9 | 27.8 | 28.2 | 28.3 | 27.6 | 28.0 | 29.6 | 28.6 | 29.5 | - |
es | SmolSent | -1.1 | 62.4 | 61.3 | 51.3 | 61.3 | 61.3 | - | - | 63.0 | - | 63.5 | 61.8 |
sat | SmolSent | -1.3 | 32.4 | 30.9 | 0.6 | 31.1 | 30.8 | 34.7 | 36.0 | 36.3 | 1.8 | 35.7 | - |
nd | Both | -1.4 | 43.9 | 41.5 | 42.5 | 42.5 | 43.2 | 42.3 | 42.9 | 44.5 | 43.6 | - | - |
qu | SmolSent | -1.9 | 34.7 | 32.8 | 29.3 | 32.8 | 33.0 | 35.3 | 35.1 | 34.0 | 22.0 | 36.3 | 27.9 |
lus | SmolSent | -2.1 | 42.6 | 39.7 | 37.2 | 40.5 | 41.4 | 40.6 | 41.5 | 43.8 | 33.5 | 42.6 | 39.0 |
scn | SmolDoc | -2.2 | 52.4 | 47.3 | 51.2 | 50.2 | 50.8 | 49.9 | 51.4 | 52.1 | 49.5 | 53.3 | 51.0 |
ne | SmolDoc | -2.7 | 54.3 | 51.8 | 51.9 | 51.6 | 52.1 | 52.4 | 52.5 | 52.4 | 52.7 | 54.9 | 45.2 |
pa-A. | SmolSent | -3.0 | 38.1 | 35.8 | 0.3 | 35.1 | 35.7 | 41.6 | 41.2 | 36.7 | 37.3 | 43.5 | - |
aeb | Both | -3.3 | 46.5 | 41.9 | 42.7 | 43.2 | 43.4 | 45.9 | 46.8 | 47.6 | 49.2 | - | 43.8 |
bo | SmolSent | -4.3 | 41.3 | 36.7 | 34.5 | 37.0 | 37.3 | 42.6 | 42.1 | 43.3 | 19.8 | 41.8 | 36.9 |
pcm | Both | -6.5 | 47.9 | 43.5 | 39.7 | 41.4 | 41.6 | 51.3 | 45.7 | 49.8 | 56.0 | - | - |
tcy | SmolDoc | -6.8 | 34.7 | 22.0 | 28.0 | 27.9 | 28.8 | 28.2 | 29.3 | 36.7 | 21.6 | 39.1 | - |
ktu | Both | -9.4 | 56.6 | 59.3 | 40.1 | 47.2 | 51.3 | 45.8 | 48.4 | 57.8 | 22.3 | 64.3 | - |
Appendix G Complete Per-Language details: the Big-Smol table
A summary of all Smol language pairs and coarse-grained information about them can be seen in Table LABEL:tab:full-details. Numbers are given in terms of examples; keep in mind that a single example in SmolDoc is a document, whereas in SmolSent it is a sentence.
Lang. pair | target language name | ISO 15924 Script | Continent | SmolSent ex | SmolDoc ex |
am_ar | Arabic | Arab | Asia | 0 | 329 |
am_zh | Mandarin Chinese | Hans | Asia | 0 | 329 |
en_aa | Afar | Latn | Africa | 863 | 66 |
en_ach | Acoli | Latn | Africa | 863 | 66 |
en_aeb | Tunisian Arabic | Arab | Africa | 862 | 66 |
en_af | Afrikaans | Latn | Africa | 863 | 130 |
en_ahr | Ahirani | Deva | Asia | 0 | 457 |
en_ak | Twi | Latn | Africa | 863 | 260 |
en_alz | Alur | Latn | Africa | 863 | 66 |
en_am | Amharic | Ethi | Africa | 863 | 584 |
en_apd | Sudanese Arabic | Arab | Africa | 855 | 66 |
en_ar-MA | Morrocan Arabic | Arab | Africa | 863 | 260 |
en_arz | Egyptian Arabic | Arab | Africa | 863 | 260 |
en_ay | Aymara | Latn | Americas | 863 | 0 |
en_ayl | Libyan Arabic | Arab | Africa | 863 | 66 |
en_bci | Baoulé | Latn | Africa | 863 | 66 |
en_bem | Bemba (Zambia) | Latn | Africa | 863 | 66 |
en_ber | Tamazight (Tifinagh Script) | Tfng | Africa | 862 | 130 |
en_ber-Latn | Tamazight (Latin Script) | Latn | Africa | 862 | 130 |
en_bfq | Badaga | Taml | Asia | 0 | 457 |
en_bfy | Bagheli | Deva | Asia | 0 | 457 |
en_bgq | Bagri | Deva | Asia | 0 | 457 |
en_bm | Bambara | Latn | Africa | 863 | 260 |
en_bns | Bundeli | Deva | Asia | 0 | 456 |
en_bo | Tibetan | Tibt | Asia | 863 | 0 |
en_bra | Braj | Deva | Asia | 0 | 457 |
en_brx | Bodo (India) | Deva | Asia | 863 | 0 |
en_ccp-Latn | Chakma (Latin script) | Latn | Asia | 0 | 457 |
en_cgg | Chiga | Latn | Africa | 863 | 66 |
en_dhd | Dhundari | Deva | Asia | 0 | 456 |
en_din | Dinka | Latn | Africa | 863 | 66 |
en_doi | Dogri | Deva | Asia | 0 | 454 |
en_dov | Dombe | Latn | Africa | 863 | 66 |
en_dyu | Dyula | Latn | Africa | 863 | 66 |
en_ee | Ewe | Latn | Africa | 863 | 130 |
en_efi | Efik | Latn | Africa | 863 | 66 |
en_ff | Fulfulde | Latn | Africa | 862 | 260 |
en_fon | Fon | Latn | Africa | 863 | 66 |
en_gaa | Ga | Latn | Africa | 863 | 66 |
en_gn | Guarani | Latn | Americas | 863 | 0 |
en_grt-Latn | Garo (Latin script) | Latn | Asia | 0 | 457 |
en_ha | Hausa | Latn | Africa | 863 | 584 |
en_hoc-Wara | Ho (Warang Chiti script) | Wara | Asia | 0 | 457 |
en_ig | Igbo | Latn | Africa | 863 | 391 |
en_kfy | Kumaoni | Deva | Asia | 0 | 457 |
en_kg | Kongo | Latn | Africa | 863 | 66 |
en_ki | Kikuyu | Latn | Africa | 863 | 66 |
en_kl | Kalaallisut | Latn | Americas | 863 | 0 |
en_kr | Kanuri | Latn | Africa | 863 | 66 |
en_kri | Krio | Latn | Africa | 863 | 130 |
en_kru | Kurukh | Deva | Asia | 0 | 457 |
en_ks | Kashmiri | Arab | Asia | 863 | 0 |
en_ks-Deva | Kashmiri (Devanagari script) | Deva | Asia | 863 | 0 |
en_ktu | Kituba (DRC) | Latn | Africa | 863 | 66 |
en_lep | Lepcha | Lepc | Asia | 0 | 456 |
en_lg | Luganda | Latn | Africa | 863 | 260 |
en_lif-Limb | Limbu (Limbu script) | Limb | Asia | 0 | 457 |
en_ln | Lingala | Latn | Africa | 863 | 260 |
en_lu | Kiluba (Luba-Katanga) | Latn | Africa | 863 | 66 |
en_luo | Luo | Latn | Africa | 863 | 260 |
en_lus | Mizo | Latn | Asia | 863 | 0 |
en_mag | Magahi | Deva | Asia | 0 | 456 |
en_mfe | Morisien | Latn | Africa | 863 | 66 |
en_mg | Malagasy | Latn | Africa | 863 | 391 |
en_mjl | Mandeali | Deva | Asia | 0 | 457 |
en_mni-Mtei | Meiteilon (Manipuri) | Mtei | Asia | 863 | 0 |
en_mos | Mossi | Latn | Africa | 863 | 66 |
en_mtr | Mewari | Deva | Asia | 0 | 457 |
en_nd | North Ndebele | Latn | Africa | 863 | 66 |
en_ndc-ZW | Ndau | Latn | Africa | 863 | 66 |
en_ne | Nepali | Deva | Asia | 0 | 456 |
en_noe | Nimadi | Deva | Asia | 0 | 315 |
en_nr | South Ndebele | Latn | Africa | 863 | 66 |
en_nso | Sepedi | Latn | Africa | 863 | 130 |
en_ny | Chichewa | Latn | Africa | 863 | 260 |
en_nqo | NKo | Nkoo | Africa | 863 | 66 |
en_om | Oromo | Latn | Africa | 863 | 391 |
en_pa-Arab | Lahnda Punjabi (Pakistan) | Arab | Asia | 863 | 0 |
en_pcm | Nigerian Pidgin | Latn | Africa | 863 | 130 |
en_qu | Quechua | Latn | Americas | 863 | 0 |
en_rn | Rundi | Latn | Africa | 863 | 66 |
en_rw | Kinyarwanda | Latn | Africa | 863 | 260 |
en_sa | Sanskrit | Deva | Asia | 863 | 0 |
en_sat | Santali (Ol Chiki script) | Olck | Asia | 863 | 0 |
en_sat-Latn | Santali (Latin Script) | Latn | Asia | 863 | 0 |
en_scl | Shina | Arab | Asia | 0 | 457 |
en_scn | Sicilian | Latn | Europe | 0 | 100 |
en_sd-Deva | Sindhi (Devanagari script) | Deva | Asia | 0 | 456 |
en_sgj | Surgujia | Deva | Asia | 0 | 356 |
en_sjp | Surjapuri | Deva | Asia | 0 | 299 |
en_sn | Shona | Latn | Africa | 863 | 260 |
en_so | Somali | Latn | Africa | 862 | 260 |
en_spv | Sambalpuri | Orya | Asia | 0 | 457 |
en_ss | Swati | Latn | Africa | 863 | 66 |
en_st | Sesotho | Latn | Africa | 863 | 260 |
en_sus | Susu | Latn | Africa | 863 | 66 |
en_sw | Swahili | Latn | Africa | 863 | 584 |
en_tcy | Tulu | Knda | Asia | 0 | 451 |
en_ti | Tigrinya | Ethi | Africa | 863 | 260 |
en_tiv | Tiv | Latn | Africa | 863 | 66 |
en_tn | Tswana | Latn | Africa | 863 | 66 |
en_trp | Kok Borok | Latn | Asia | 0 | 457 |
en_ts | Tsonga | Latn | Africa | 863 | 66 |
en_tum | Tumbuka | Latn | Africa | 863 | 66 |
en_unr-Deva | Mundari (Devanagari script) | Deva | Asia | 0 | 457 |
en_ve | Venda | Latn | Africa | 863 | 66 |
en_wbr | Wagdi | Deva | Asia | 0 | 455 |
en_wo | Wolof | Latn | Africa | 863 | 260 |
en_xh | Xhosa | Latn | Africa | 863 | 260 |
en_xnr | Kangri | Deva | Asia | 0 | 457 |
en_xsr-Tibt | Sherpa (Tibetan script) | Tibt | Asia | 0 | 457 |
en_yo | Yoruba | Latn | Africa | 863 | 584 |
en_yue | Cantonese | Hant | Asia | 863 | 584 |
en_zu | Zulu | Latn | Africa | 863 | 260 |
sw_ar | Arabic | Arab | Asia | 0 | 330 |
sw_zh | Mandarin Chinese | Hans | Asia | 0 | 330 |

### v2

@@ Appendix G Full results
| gaa | Both | +23.1 | 22.7 | 44.5 | 44.4 | 45.8 | 47.4 | 34.7 | 44.0 | 40.9 | 6.6 | 48.3 | - |
| dov | Both | +21.1 | 19.1 | 39.2 | 38.3 | 40.2 | 40.6 | 19.2 | 39.5 | 18.2 | 8.7 | 41.7 | - |
| ahr | neither | +17.8 | 24.2 | 31.8 | 41.9 | 42.0 | 42.8 | 32.8 | 39.0 | 30.0 | 36.9 | - | - |
| sus | Both | +17.8 | 11.3 | 28.3 | 26.8 | 29.1 | 30.3 | 26.1 | 29.4 | 20.7 | 5.6 | 34.6 | - |
| nqo | Both | +17.5 | 0.2 | 17.9 | 17.1 | 17.7 | 17.5 | 17.1 | 17.9 | 17.2 | 1.1 | 19.1 | - |
| alz | Both | +15.5 | 16.9 | 31.5 | 30.5 | 32.4 | 33.4 | 25.3 | 30.6 | 26.9 | 8.9 | 36.6 | - |
| lu | Both | +13.8 | 27.6 | 37.5 | 39.3 | 41.4 | 42.2 | 27.2 | 37.0 | 34.8 | 21.9 | - | - |
@@
| ks | neither | +7.3 | 22.1 | 29.4 | 0.4 | 29.4 | 29.7 | 28.0 | 30.4 | 30.5 | 26.3 | - | 36.7 |
| nr | neither | +7.0 | 48.9 | 54.0 | 51.1 | 55.9 | 57.5 | 48.0 | 53.6 | 51.2 | 45.5 | 59.5 | - |
| doi | neither | +6.6 | 34.3 | 28.1 | 41.4 | 40.9 | 41.3 | 35.9 | 39.5 | 38.2 | 27.7 | 40.4 | - |
| sat-L. | SmolSent | +6.4 | 12.8 | 19.5 | 15.5 | 19.2 | 20.8 | 25.3 | 22.5 | 22.7 | 21.3 | 22.7 | - |
| mfe | neither | +5.3 | 59.5 | 65.4 | 62.6 | 64.8 | 66.9 | 59.6 | 65.0 | 59.8 | 59.5 | 67.5 | - |
| ach | Both | +5.2 | 33.2 | 43.1 | 32.5 | 38.4 | 39.2 | 32.4 | 37.3 | 35.1 | 23.8 | 43.2 | - |
| ayl | neither | +4.5 | 47.3 | 51.7 | 51.9 | 51.8 | 53.9 | 45.3 | 48.9 | 46.3 | 48.6 | - | - |
| st | neither | +4.5 | 49.9 | 54.0 | 55.2 | 54.4 | 55.0 | 49.4 | 57.0 | 53.1 | 49.2 | 49.0 | 47.2 |
| ber-L. | Both | +4.2 | 26.1 | 27.9 | 30.3 | 30.3 | 30.9 | 27.6 | 32.7 | 32.1 | 21.2 | 34.7 | - |
| apd-S. | neither | +3.6 | 42.3 | 50.2 | 43.3 | 45.9 | 47.1 | 43.2 | 45.0 | 42.5 | 45.6 | - | - |
| ve | neither | +3.5 | 47.9 | 50.0 | 48.7 | 51.4 | 52.2 | 50.2 | 53.1 | 52.7 | 43.9 | 56.8 | - |
| kri | neither | +2.8 | 31.5 | 34.2 | 31.8 | 34.3 | 34.7 | 34.5 | 33.5 | 30.7 | 34.9 | 34.9 | - |
| tiv | Both | +2.5 | 23.8 | 25.7 | 25.8 | 26.3 | 26.5 | 22.3 | 24.2 | 24.5 | 1.5 | 25.2 | - |
| gn | SmolSent | +2.3 | 37.4 | 37.8 | 30.4 | 39.7 | 38.4 | 36.0 | 38.0 | 36.4 | 35.6 | 38.4 | 38.5 |
| mos | Both | +2.3 | 18.2 | 20.9 | 18.9 | 20.5 | 21.1 | 24.3 | 25.0 | 20.9 | 1.3 | - | 23.8 |
| tum | neither | +1.1 | 40.8 | 39.5 | 42.4 | 41.9 | 42.8 | 40.0 | 42.7 | 43.8 | 37.7 | 45.4 | 36.2 |
| ti | neither | +0.8 | 24.2 | 24.3 | 24.9 | 25.0 | 25.7 | 25.0 | 26.2 | 26.1 | 9.3 | 26.1 | 25.5 |
| yo | neither | +0.6 | 34.6 | 33.8 | 35.4 | 35.2 | 35.8 | 29.2 | 36.8 | 26.7 | 27.6 | 21.3 | 32.6 |
| tn | neither | +0.2 | 52.5 | 50.8 | 51.9 | 52.7 | 53.2 | 50.1 | 51.7 | 53.3 | 36.8 | 55.6 | 53.0 |
| ar-M. | neither | +0.1 | 40.1 | 41.8 | 39.1 | 40.2 | 40.9 | 40.5 | 40.8 | 40.4 | 41.0 | - | 43.0 |
| am | neither | +0.1 | 34.0 | 33.2 | 33.0 | 34.1 | 33.5 | 31.6 | 32.6 | 35.8 | 29.6 | 34.7 | 30.3 |
| ig | neither | -0.1 | 47.2 | 47.1 | 47.0 | 47.1 | 47.8 | 43.9 | 46.2 | 47.8 | 46.2 | 47.6 | 46.6 |
| so | neither | -0.1 | 49.7 | 46.2 | 50.0 | 49.6 | 49.1 | 48.7 | 49.8 | 50.3 | 50.8 | 50.6 | 48.6 |
| arz | neither | -0.1 | 48.6 | 46.1 | 48.9 | 48.5 | 47.8 | 49.6 | 50.3 | 48.8 | 49.7 | - | 49.6 |
| kl | SmolSent | -0.3 | 40.6 | 39.7 | 30.2 | 40.3 | 41.2 | 42.2 | 43.1 | 41.2 | 41.6 | 42.9 | - |
| sa | SmolSent | -0.3 | 33.0 | 32.9 | 26.7 | 32.7 | 33.4 | 31.8 | 33.0 | 32.1 | 32.0 | 35.2 | 29.0 |
| ay | SmolSent | -0.5 | 32.7 | 31.8 | 24.2 | 32.2 | 32.4 | 33.4 | 33.2 | 32.9 | 30.0 | 34.7 | 31.7 |
| sn | neither | -0.5 | 50.5 | 48.3 | 50.2 | 50.0 | 50.3 | 46.8 | 48.8 | 51.8 | 50.3 | 49.2 | 48.2 |
@@
| ss | neither | -0.6 | 50.2 | 48.8 | 48.2 | 49.6 | 50.3 | 49.6 | 51.2 | 51.8 | 46.0 | 56.3 | 48.1 |
| yue | neither | -0.7 | 26.8 | 25.8 | 25.1 | 26.1 | 26.1 | 28.2 | 28.2 | 27.5 | 31.6 | 25.9 | 22.6 |
| bci | Both | -0.7 | 23.2 | 22.2 | 21.8 | 22.5 | 22.9 | 17.1 | 20.7 | 27.6 | 1.0 | 29.8 | - |
| ndc-Z. | neither | -1.0 | 29.2 | 27.9 | 28.9 | 28.2 | 28.3 | 27.6 | 28.0 | 29.6 | 28.6 | 29.5 | - |
| es | neither | -1.1 | 62.4 | 61.3 | 51.6 | 61.3 | 61.3 | - | - | 63.0 | - | 63.5 | 61.8 |
| sat | SmolSent | -1.3 | 32.4 | 30.9 | 1.0 | 31.1 | 30.8 | 34.7 | 36.0 | 36.3 | 1.8 | 35.7 | - |
| rw | neither | -1.4 | 45.2 | 43.1 | 43.0 | 43.8 | 43.8 | 43.2 | 44.0 | 45.1 | 44.7 | 48.8 | 43.4 |
| nd | neither | -1.4 | 43.9 | 41.5 | 42.6 | 42.5 | 43.2 | 42.3 | 42.9 | 44.5 | 43.6 | - | - |
| sw | neither | -1.5 | 66.7 | 64.6 | 64.2 | 65.2 | 64.8 | 64.0 | 65.5 | 67.2 | 66.5 | 65.3 | 60.5 |
| mg | neither | -1.9 | 52.8 | 48.9 | 51.6 | 50.9 | 51.5 | 52.4 | 52.5 | 53.3 | 52.2 | 52.6 | 52.1 |
| qu | SmolSent | -1.9 | 34.7 | 32.8 | 30.4 | 32.8 | 33.0 | 35.3 | 35.1 | 34.0 | 22.0 | 36.3 | 27.9 |
| zu | neither | -2.0 | 58.3 | 56.4 | 55.3 | 56.3 | 56.1 | 54.1 | 55.5 | 58.5 | 57.5 | 57.6 | 57.6 |
| lus | SmolSent | -2.1 | 42.6 | 39.7 | 38.0 | 40.5 | 41.4 | 40.6 | 41.5 | 43.8 | 33.5 | 42.6 | 39.0 |
| scn | neither | -2.2 | 52.4 | 47.3 | 50.0 | 50.2 | 50.8 | 49.9 | 51.4 | 52.1 | 49.5 | 53.3 | 51.0 |
| nso | neither | -2.3 | 46.8 | 42.8 | 43.6 | 44.5 | 44.6 | 46.9 | 47.7 | 48.1 | 46.9 | 47.6 | 45.5 |
| xh | neither | -2.3 | 53.9 | 49.7 | 50.7 | 51.6 | 51.8 | 51.6 | 52.3 | 53.9 | 53.7 | 54.8 | 51.2 |
| ne | neither | -2.7 | 54.3 | 51.8 | 51.7 | 51.6 | 52.1 | 52.4 | 52.5 | 52.4 | 52.7 | 54.9 | 45.2 |
| pa-A. | neither | -3.0 | 38.1 | 35.8 | 0.3 | 35.1 | 35.7 | 41.6 | 41.2 | 36.7 | 37.3 | 43.5 | - |
| aeb | neither | -3.3 | 46.5 | 41.9 | 42.3 | 43.2 | 43.4 | 45.9 | 46.8 | 47.6 | 49.2 | - | 43.8 |
| ha | neither | -3.4 | 54.5 | 50.7 | 49.9 | 51.1 | 51.5 | 50.9 | 51.0 | 54.1 | 53.9 | 53.8 | 53.9 |
| ts | neither | -3.6 | 50.6 | 47.2 | 46.4 | 47.0 | 48.1 | 49.7 | 50.1 | 51.6 | 49.0 | 52.9 | 51.3 |
| rn | neither | -3.9 | 44.5 | 39.3 | 40.5 | 40.6 | 40.9 | 43.1 | 43.4 | 46.2 | 44.3 | 45.4 | 45.0 |
| af | neither | -4.2 | 71.9 | 68.8 | 67.9 | 67.7 | 68.3 | 71.7 | 72.5 | 72.1 | 71.8 | 71.5 | 68.6 |
| bo | SmolSent | -4.3 | 41.3 | 36.7 | 34.7 | 37.0 | 37.3 | 42.6 | 42.1 | 43.3 | 19.8 | 41.8 | 36.9 |
| ny | neither | -5.1 | 55.0 | 47.5 | 50.5 | 49.9 | 49.7 | 53.0 | 53.1 | 55.3 | 53.9 | 55.8 | 50.3 |
| pcm | neither | -6.5 | 47.9 | 43.5 | 39.4 | 41.4 | 41.6 | 51.3 | 45.7 | 49.8 | 56.0 | - | - |
| tcy | SmolDoc | -6.8 | 34.7 | 22.0 | 28.1 | 27.9 | 28.8 | 28.2 | 29.3 | 36.7 | 21.6 | 39.1 | - |
| ktu | Both | -9.4 | 56.6 | 59.3 | 40.4 | 47.2 | 51.3 | 45.8 | 48.4 | 57.8 | 22.3 | 64.3 | - |